Calculate 25°C in degree Fahrenheit. Report to 2 significant figures
Answer #1

Formula :

C/5 = F- 32 /9

C = 25 ° C

25 / 5 = F - 32 / 9

5 = F-32 / 9

4×9 = F-32

45 + 32 = F

F = 77 ° F